OpenStudy (anonymous):

integrate root (secx+1)

OpenStudy (experimentx):

there seems no closed expression for this integration.

OpenStudy (anonymous):

What about this: \[\int\sqrt{\sec x+1}\;dx\cdot\frac{\sqrt{\sec x-1}}{\sqrt{\sec x-1}}\\ \int\frac{\sqrt{\sec^2 x-1}}{\sqrt{\sec x-1}}\;dx\\ \int\frac{\tan x}{\sqrt{\sec x-1}}\;dx\\ \int\frac{\tan x}{\sqrt{\sec x-1}}\;dx\cdot\frac{\sec x}{\sec x}\\ \int\frac{\sec x\tan x}{\sec x\sqrt{\sec x-1}}\;dx\] \[u=\sec x\\ du=\sec x\tan x\;dx\] \[\int\frac{1}{u\sqrt{u-1}}\;du\] Still no closed form? I'm not sure.

OpenStudy (experimentx):

looks like wolfram was giving me wrong result, Here is result from Maple

OpenStudy (experimentx):

looks like the result from the maple is correct ... differentiating and simplifying it yields \[ \frac{(1+cos(x))}{(sqrt((1+cos(x))/cos(x))*cos(x))} \] which is equal to the integrand.

OpenStudy (experimentx):

looks like I shouldn't rely too much on software ... using the result from above by @SithsAndGiggles yields \[ \int \sqrt{1 + \sec x} dx = 2 \tan^{-1} \left( \sec x - 1\right) + c\] differentiating which you get the result, your integrand. But the problem is \[ \tan x = \sqrt{ \sec^2x - 1 }\] is a conditional expression where tan is positive. I am not sure if it should be the result.